The ratio between the ages of two pwrsons is 6:5 and sum of there ages is 77 then how many years later there ratio becomes 8:7
Answer by mananth(16946)

let the ages be x & y
x/y = 6/5
x=6y/5
...
x+y =77
plug vaue of x in the equation
6y/5 + y = 77
6y+5y / 5 = 77
11y= 5*77
y = 5*77/11
y = 35
x=42
..
let x be the number of years after which their ratios become 8/7
42+x/35+x = 8/7
8(35+x)=7(42+x)
280+8x=294+7x
8x-7x=294-280
x= 14 .
Ater 14 years their ratio will be 8/7
...
m.ananth@hotmail.ca